Nasa is expected to send a 2600 kg satellite 450 km above the earth surface.


A) what is its radius?


B) what speed will it have?( Hint: earth mass is 5.97 x 10²⁴ kg)


C) what is its orbital period?


D) what is its radial acceleration?

1
Expert's answer
2022-01-07T09:30:04-0500

A) The radius of the orbit (Earth's radius is 6370 km):


"R=R_E+H=6.82·10^6\text{ m}."


B) The speed:


"v=\sqrt\dfrac{GM_E}{R}=7643\text{ m/s}."


C) The orbital period:


"T=\dfrac{2\pi R}{v}=1\text{ h 33 min 27 s}."


D) The centripetal acceleration:


"a_c=\dfrac{v^2}{R}=8.57\text{ m/s}^2."